# The representation of Convectively Coupled Equatorial Waves and upscale energy transfer in models with explicit and parametrized convection

### v1

@@ Abstract
Convectively Coupled Equatorial Waves (CCEWs) dominate atmospheric variability on timescales of 2–30 days in the Tropics, bringing episodes of widespread heavy precipitation. This study compares the representation of CCEWs and their connection to upscale energy transfer in two Met Office Unified Model simulations of the full tropical channel with identical km-scale resolution. The principal difference between the simulations is that one parametrizes convection (GAL), while the other (RAL) is convection permitting. This means GAL acts to remove vertical instability without explicitly representing the resolved-scale circulation associated with convective plumes. We present the first quantitative diagnosis of interscale energy transfer and its relation to CCEWs. This diagnosis is important because upscale energy transfer between convection and large-scale waves may influence accurate simulation and predictability of tropical weather systems. The average upper-tropospheric upscale transfer simulated by RAL is approximately 50% higher than GAL. CCEWs are more coherent in RAL, with an average phase-speed variability 80% higher than observations, compared with 166% higher in GAL. RAL also simulates greater upscale energy transfer within waves than GAL with a stronger correlation between the interscale energy transfer rate and equatorial wave winds. Simulated Kelvin and Rossby waves are associated with upscale energy transfer from scales 2-8 times smaller than the dominant wavelength, related to active deep convection within a particular sector of the wave phase. Our findings show that the explicit representation of convective motions have a significant impact on the simulation of upscale energy transfer, and are therefore very likely to be a significant factor in the faithful simulation of the convective coupling within CCEWs.
@@ 1 Introduction
There also remains the outstanding question of whether forecast errors propagate upscale or up-amplitude. A comprehensive review of this question is beyond the scope of this paper, but the authors point to the recent work of Rotunno et al. (2023), where it is reported that one must be careful with the methods of reducing to spectra as to whether you are examining upscale or up-amplitude error growth. In the case of operational ensembles, small-amplitude initial condition perturbations are applied at all scales so error growth in these systems may be termed “up-amplitude”. We are examining a deterministic system with no prescribed initial perturbations and explicitly examining local spectra, so the energy transfers we describe are upscale.
@@ 2 Local interscale energy transfers
$\mathcal{D}_{L}(\mathbf{u})=\frac{1}{4}\int\mathrm{d}\mathbf{r}(\nabla G_{L})(\mathbf{r})\cdot\delta\mathbf{u}(\mathbf{r})|\delta\mathbf{u}(\mathbf{r})|^{2}.$ (3)
Fundamental turbulence research typically considers unstratified 3D flows far away from boundaries; hence spherically-symmetric filter kernels are usually employed. However, in the geosciences, the importance of gravity singles out the vertical direction as special. Therefore we consider only 2D filter kernels acting on surfaces perpendicular to gravity. We assume that these surfaces are concentric to the Earth’s surface, which will itself be assumed to be spherical. For a sphere of radius $R$ (denoted $S^{2}(R)$), the surface element $\mathrm{d}S$ may be written:
$\mathrm{d}S=R\sin\alpha\mathrm{d}\chi R\mathrm{d}\alpha=R\sin\left(\frac{r}{R}\right)\mathrm{d}\chi\mathrm{d}r,$
where $r$ is the great circle distance from $\mathbf{x}$, $\alpha$ is the angle subtended by the surface of the sphere between $\mathbf{x}$ and $r$, and $\chi$ the angle between the arc of constant latitude and $\mathbf{r}$. For $\frac{r}{R}\ll 1$ this reduces to $\mathrm{d}S\approx r\mathrm{d}\chi\mathrm{d}r$, matching the surface element of Euclidean space expressed in polar coordinates.
For any 2D kernel that depends only on the magnitude of $\mathbf{r}$, integrating over a spherical surface, (3) becomes:
$\int_{S^{2}(R)}\nabla G_{L}(\mathbf{r})\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}\mathrm{d}\mathbf{r}$ $=\int_{0}^{\pi R}\int_{0}^{2\pi}\frac{\partial G_{L}(r)}{\partial r}\hat{\mathbf{r}}\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}R\sin\left(\frac{r}{R}\right)\mathrm{d}\phi\ \mathrm{d}r$
$=R\int_{0}^{\pi R}\frac{\partial G_{L}(r)}{\partial r}\sin\left(\frac{r}{R}\right)\left[\int_{0}^{2\pi}\hat{\mathbf{r}}\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}\mathrm{d}\phi\right]\mathrm{d}r.$ (4)
Since the sphere is compact, it only really makes sense to consider compactly-supported filter kernels. One such compactly supported filter kernel is the “standard mollifier” used by Kuzzay et al. (2015, 2017); Faranda et al. (2018), defined by:
@@
To compute the $\mathcal{D}_{L}(\mathbf{u})$ field used in this study, we first calculate velocity increments from the $(u,\upsilon,w)$ field. We note here the inclusion of vertical velocity $w$ – reported to be of negligible importance in the oceanic analysis done by Storer et al. (2023), it is an important variable in the tropical atmosphere where there is abundant deep convection. Then, an angular average (around a circle of radius $|\mathbf{r}|$) is applied at each point. This is the quantity in square brackets in (4) Our method is also less sensitive to noise because gradients are computed analytically on smooth filter kernels rather than on discrete model fields as is done in other methods (e.g. Aluie et al. (2018)). The result of the angular integral is then integrated over $\mathbf{r}$, weighted by the gradient of the filter kernel. The method outlined here is fully described in Shipley et al. (in prep.).
To give an idea of what the interscale energy transfer field looks like in the tropical atmosphere, an instantaneous snapshot of $\mathcal{D}_{L}(\mathbf{u})$ from RAL for $L=440$ km at 200 hPa is shown in Fig. 1. The large mesoscale of 440 km is used for the majority of figures in this paper, and the energy transfer across this length scale is calculated from N2560 model data interpolated onto a N1280 or 0.5 degree grid(depending on length scale) to reduce computational expense. Figure 1 shows that this choice is justified, since at such a large length scale, the calculation whether made on the original N2560 grid, a grid half as fine or the 0.5 degree grid makes little difference. The 440 km length scale is chosen to facilitate comparison with Faranda et al. (2018), who elected to study energy transfers across 220 km. The reason our chosen $L$ is twice that of Faranda et al. (2018)’s is because the kernel we use is the same as theirs, but we find it more informative to talk about length scales on the diameter of the kernel rather than the radius. Furthermore, this length scale is well separated from small scale convection and the wavelength of CCEWs. The domain in Fig. 1 (15S to 20N) is chosen to highlight the activity north of the equator during DYAMOND Summer (boreal summer - the ITCZ is centred north of the equator during this season).
@@ 3.1 The K-Scale simulations
The CTC km-scale model domain is being run as part of the UK Met Office’s ‘K-Scale’ initiative (Jones et al. (2023), Jones et al. (In review)) which also serves as the UK contribution to the DYAMOND protocol (Stevens et al., 2019a; Takasuka et al., 2024). The DYAMOND (DYnamics of the Atmosphere Modelled On Non-hydrostatic Domains) protocol is an international collaborative project involving modelling centres across the world exploring CPMs spanning global or large domains at high resolution and their benefits. All centres modelled a common time period. We will focus on the DYAMOND summer period in this study from 2016-08-01 to 2016-09-09. The DYAMOND Summer period was chosen to match up with the NARVAL (Stevens et al., 2019b) observational campaign during these dates. 40 days was chosen to provide a month of data after any spin up and be long enough to separate questions of predictability and internal variability. The relevant tropical climate indices during DYAMOND Summer are listed in Table 1. Of particular relevance to this study is that there is no MJO during DYAMOND Summer, meaning CCEWs can be studied as their own system rather than as part of the MJO envelope (e.g. Roundy (2008); Kiladis et al. (2009)). The output data used are 3-hourly. The simulations are free-running for 40 days, i.e. with no data assimilation scheme in place like for an operational forecast. This is to allow features to grow by themselves in the model rather than be forced by the initial conditions fed in from the data assimilation scheme.
Table 1: Climate indices during the DYAMOND Summer study period. Abbreviations are the El Niño Southern Oscillation (ENSO), Indian Ocean Dipole (IOD), Madden-Julian Oscillation (MJO), and Quasi-Biennial Oscillation (QBO). The indices were calculated for ENSO using the Nino3.4 index from HadISST data (Rayner et al., 2003); the IOD using the Dipole Mode Index from HadISST data; the MJO using the Realtime Multivariate MJO index from NOAA OLR and NCEP/NCAR reanalysis winds data (Wheeler and Hendon, 2004; Kalnay et al., 1996), and the QBO from NCEP/NCAR reanalysis data (Kalnay et al., 1996).
@@
ENSO Neutral to weak La Niña
IOD Strong negative phase
@@
QBO Westerly
For the analysis carried out in this study, we compare two UM (Brown et al., 2012) coupled to the Joint UK Land Environment Simulator (JULES, Best et al. 2011; Clark et al. 2011) simulations from the K-Scale hierarchy (Jones et al., In review) with differing science configurations. One simulation (referred to as GAL9 N2560 or GAL for short throughout) makes use of the Global Atmosphere and Land version 9 (GAL9, Walters et al. 2019) configuration. The other simulation (RAL3 N2560, RAL for short) uses the Regional Atmosphere and Land version 3.2 (RAL3.2, Bush et al. 2025). Details of the differences between the two simulations are summarised in Table 2.
Both models have identical grids and cover the CTC domain from 40S to 26N. These bounds were chosen to meet the Himalayas in the north and encompass all of southern Africa and the Indian Ocean tropical cyclone basin. The two models are both driven by a global N1280 ($\Delta x\sim 10$ km) model with GAL9 physics. The key difference between GAL and RAL being investigated in this study is the convection scheme (or lack thereof), with most other features being kept the same.
GAL9 employs a mass flux parametrization of shallow and mid-level convection. The convection parametrization is not active in RAL3.2, so any convection occurring is permitted to explicitly evolve through the model dynamics on the model grid. There is additional sub-grid vertical mixing in RAL3.2 above the boundary layer that is parametrized by a Smagorinsky-Lilly-type eddy diffusivity. Other differences between RAL and GAL that must be notes are the different cloud fraction schemes: GAL9 uses the PC2 prognostic scheme (Wilson et al., 2008), and RAL3.2 the bimodal diagnostic scheme (Van Weverberg et al., 2021a, b). The microphysics in GAL9 is parametrized using the Wilson and Ballard (1999) single moment scheme; and RAL3.2 uses the double moment Cloud AeroSol Interacting Microsphysics scheme (CASIM, Field et al. 2023). In the boundary layer, GAL9 employs the 1D turbulent mixing scheme of Lock et al. (2000), and RAL3 employs a blended 3D Smagorinsky and 1D scheme, described in Boutle et al. (2014).
@@ 3.2 Equatorial wave filtering
Equatorial wave dynamical fields are extracted from the model or analysis $(u,\upsilon,Z)$ (zonal wind, meridional wind, geopotential height) fields using the 2D-in-Space Parabolic Cylinder Function (2DS-PCF) method developed by Yang et al. (2003). Due to the relatively short time period covered by our simulations, the “realtime” extension of the method (Yang et al., 2021) is used, whereby a long analysis period is appended to the start of the forecast to enable a time filter with a very broad window so that eastward and westward signals can be distinguished before spatial projection. There are multiple ways to identify equatorial waves from modelled and observed flows, with Knippertz et al. (2022) calling for a combination of wavenumber-frequency filtering and spatial projection methods. Since we are concerned here with kinetic energy and the 3D dynamical structure of the waves, the 2DS-PCF method is the most suitable for our purposes.
@@ 4.1 Simulated waves and precipitation
Hovmöller (longitude-time) diagrams of quantities averaged across 10S-10N are shown in Fig. 2 - the first 20 days are chosen for clarity in the figure, and to allow us to analyse the evolution of simulated features from the same initial conditions. This diagram gives us a first look at the key differences between the RAL and GAL models and the observations. In terms of precipitation: the GAL model has many more stationary features than the RAL or observations (e.g. see the features around 90W at the start of the simulation), and its eastward propagating features are quite weak. This key difference between explicit and parametrized convection in simulating propagating precipitation is also found in the study of Dias et al. (2018). The abundance of propagating precipitation in the RAL model is more consistent with the observations than the GAL, but appears to simulate too much rain (see for example in the Indian Ocean region around 90E). These features of the models compared to the observations are also reflected in Table 3, where GAL simulates a higher mean precipitation compared to GPM, albeit with a smaller variance. RAL simulates similar variance to the observations, but with a mean that is higher still. All differences between the datasets are significantly different at the 99% confidence level due to their large size. These differences in precipitation mean and variances between the km-scale models and observations are consistent with the findings of Judt and Rios-Berrios (2021), where they analysed the DYAMOND Summer runs from the MPAS-A model.
@@
Where the wave contours travel with the precipitation, this is a clear signal of a Convectively Coupled Equatorial Wave (CCEW). The Convectively Coupled Kelvin Wave (CCKW) initialised on 2nd August at 90E and propagating across the Maritime Continent is present in both models in Figs. 2 (b) and (c) as well as in the analysis/observations in Fig. 2 (a). In the observations, this wave propagates eastward until it encounters a convectively coupled Rossby (R1) wave and WMRG over the Maritime Continent at about 120E, resulting in a large amount of precipitation being deposited over the region, and the slowing of the CCKW. This interaction does not occur in the same way in the models in (b) and (c), with there being much weaker R1 and WMRG waves in the models; in RAL the KW slows much like in the observations, but the deposition of large amounts of rain is difficult to pick out amongst the very abundant rain in that model, and in the GAL the CCKW is terminated altogether. This interaction between the Kelvin, R1 and WMRG waves does not occur in the same way in the two models because the simulated propagation speeds of the waves in RAL and GAL are different (see Table 4). In GAL, the CCKW propagates a similar distance making it as far as the dateline in the central Pacific by Aug 11th, where it also encounters a Rossby wave, albeit in the opposite phase compared to the observations. In the GAL model at this point there is also not a strong WMRG like there is in the analysis. In the RAL, the CCKW propagates farther, continuing after interaction with the easterly-phase R1 wave and terminating 180 degrees and 2 weeks later at 90W, presumably encountering the Andes. This difference may however be a factor of the free-running simulations having diverged from the analysis by 10 days out.
@@
The dominant phase speeds of the three main EW types for the UM analysis and the two models are shown in Table 4. The observed phase speeds from the analysis were cross-referenced with the extremely comprehensive analyses of Yang et al. (2007b, 2009) where they used 15 years of reanalysis data and satellite-observed convective fields, and are comparable. The models simulate a faster Kelvin wave, but the variability ranges overlap. This is the case for all the models and waves. It is also interesting to note that the models always simulate a higher standard deviation, particularly the GAL. This could be because of a reduced degree of convective coupling in GAL, allowing for a wider range of ‘dry’ (uncoupled with convection) propagating waves. Convective coupling reduces phase speed in CCEWS (Kiladis et al., 2009). For the R1 wave, the GAL simulates the same dominant phase speed with a higher variability, and the RAL a much faster phase speed with a much larger variability. The reason for this is because the distribution of phase speeds across the record calculated from the Radon transform (not shown) is bimodal. In the analysis these peaks are relatively close together at around 9 and 13 m s-1. In the RAL, the faster peak is strongest at the 15 m s-1 value reported in the table, and the second, weaker peak, is at around 6 m s-1, hence the large variability range.
@@
The differences in the simulation of the 3 key wave types can be illustrated by a longitude-height plot of standard deviation in time of wave velocity fields as shown in Fig. 3. The six-week average structures in the UM analysis presented here are similar to the 15-year average structures presented in Yang et al. (2007a, 2009). As previously reported by these earlier studies, we expect to see the strongest wave activity in the upper-troposphere for the westward waves and around the tropopause for the Kelvin waves. This is due to the the highest wave amplitude being at these levels. The peak in Kelvin wave variability around 90E at upper-tropospheric levels in the analysis is not fully simulated by either model. It is slightly weaker and further east in RAL (c), and not well defined in GAL (b). There is another peak west of 90W around 200 hPa - this is likely associated with Kelvin waves propagating across the Pacific and terminating at the Andes. This peak is stronger than analysis in GAL, and not present in RAL. The reason for this being stronger in GAL and not present in RAL could be due to the presence of the gravity wave drag (GWD) parametrization in the GAL simulation. This parametrization simulates the subgrid response to the flow blocking by unresolved orography. The N2560 resolution is in the ‘grey zone’ of the orographic GWD (some orography will be resolved, but only a fraction), meaning it is possible that the GWD is simulating too much drag in GAL, and without the parametrization in RAL, not enough is simulated. Further analysis of this particular parametrization is beyond the scope of this study; see Vosper et al. (2016) for details.
@@
The R1 waves in the analysis in Fig. 3 (d) exhibit activity centred around 200 hPa in the western hemisphere (WH) - a pattern also seen in reanalysis in Yang et al. (2009). The GAL model simulation (e) has a less pronounced peak with wave activity spread over a wider longitudinal range, while the RAL (f) has an activity centre that is also weaker but more concentrated, in this case around 0. The middle and lower troposphere structure of the R1 wave in the analysis consists of a peak in activity from 90E and eastwards. For the WMRG waves in Fig. 3 (g), there are peaks in activity centred around 200 hPa mostly in the WH (also seen in the analysis product in Yang et al. (2009)). GAL (h) represents this quite similarly, whereas RAL (i) simulates a very strong peak in activity just east of 0, and the peaks in the WH are weaker than in the analysis and GAL. The differences between GAL and RAL may just be due to noise given the short record analysed, but importantly both are able to represent the general behaviour of variability in time throughout the tropical atmosphere, shown by their similarity to the UM analysis which in turn is qualitatively similar to the patterns displayed in Yang et al. (2009).

### v2

@@ Abstract
Convectively Coupled Equatorial Waves (CCEWs) dominate atmospheric variability on timescales of 2–30 days in the Tropics, bringing episodes of widespread heavy precipitation. This study compares the representation of CCEWs and their connection to upscale energy transfer in two Met Office Unified Model simulations of the full tropical channel with identical km-scale resolution for the DYAMOND Summer period. The principal difference between the simulations is that one parametrizes convection (GAL9), while the other (RAL3) is convection permitting. The GAL9 convection scheme acts to remove vertical instability without explicitly representing the resolved-scale circulation associated with convective plumes. We present the first quantitative diagnosis of interscale energy transfer and its relation to CCEWs. This diagnosis is important because upscale energy transfer between convection and large-scale waves may influence accurate simulation of tropical weather systems. The average upper-tropospheric upscale transfer simulated by RAL3 is approximately 50% higher than GAL9. CCEWs are more coherent in RAL3, with an average phase-speed variability 80% higher than observations, compared with 166% higher in GAL. RAL3 also simulates greater upscale energy transfer within waves than GAL9 with a stronger correlation between the interscale energy transfer and equatorial wave winds. Kelvin and Rossby waves are associated with upscale energy transfer from scales 4-8 times smaller than their wavelength, related to active deep convection within a particular sector of the wave phase. Our findings show that the explicit representation of convection has a significant impact on the simulation of upscale energy transfer, and is very likely to be a significant factor in the faithful simulation of convective coupling within CCEWs.
@@ 2 Local interscale energy transfers
$\mathcal{D}_{L}(\mathbf{u})=\frac{1}{4}\int\mathrm{d}\mathbf{r}(\nabla G_{L})(\mathbf{r})\cdot\delta\mathbf{u}(\mathbf{r})|\delta\mathbf{u}(\mathbf{r})|^{2},$ (3)
where $\mathbf{x}$ is a position vector and $\mathbf{r}$ is a displacement vector.
Fundamental turbulence research typically considers unstratified 3D flows in Euclidean space far away from boundaries; hence spherically-symmetric filter kernels are usually employed. However, in the geosciences, the importance of gravity singles out the vertical direction as special. Therefore we consider only 2D filter kernels acting on surfaces perpendicular to gravity. We assume that these surfaces are concentric to the Earth’s surface, which will itself be assumed to be spherical. On a curved surface $S$, the local interscale energy transfer term is written:
$\mathcal{D}_{L}(\mathbf{u})=\frac{1}{4}\int_{S}\mathrm{d}S\ \left[\nabla G_{L}(r,\chi)\right]\cdot\delta\mathbf{u}(r,\chi)|\delta\mathbf{u}(r,\chi)|^{2},$ (4)
where $\delta\mathbf{u}(r,\chi):=\mathbf{u}(x_{2})-\mathbf{u}(x_{1})$, with $x_{1}$, $x_{2}$ two points on the surface $S$ separated by a geodesic distance $r$ at an angle $\chi\in[0,2\pi]$ from an arbitrary reference direction. For a sphere of radius $R$ (denoted $S^{2}(R)$), the surface element $\mathrm{d}S$ may be written:
$\mathrm{d}S=R\sin\alpha\mathrm{d}\chi R\mathrm{d}\alpha=R\sin\left(\frac{r}{R}\right)\mathrm{d}\chi\mathrm{d}r,$ (5)
where $r$ is the great circle distance between $x_{1}$ and $x_{2}$, $\alpha\in(0,\pi]$ is the angle subtended by the surface of the sphere between $x_{1}$ and $x_{2}$, and $\chi$ the angle between the great circle arc connecting $x_{1}$ and $x_{2}$ and the line of constant latitude passing through $x_{1}$. For $\frac{r}{R}\ll 1$ this reduces to $\mathrm{d}S\approx r\mathrm{d}\chi\mathrm{d}r$, matching the surface element of Euclidean space expressed in polar coordinates.
For any 2D kernel that depends only on the great circle distance $r$, and not on the angle $\chi$, integrating over a spherical surface, (3) becomes:
$\int_{S^{2}(R)}\nabla G_{L}(r)\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}\mathrm{d}S$ $=\int_{0}^{\pi R}\int_{0}^{2\pi}\frac{\partial G_{L}(r)}{\partial r}\hat{\mathbf{r}}\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}R\sin\left(\frac{r}{R}\right)\mathrm{d}\chi\ \mathrm{d}r$
$=R\int_{0}^{\pi R}\frac{\partial G_{L}(r)}{\partial r}\sin\left(\frac{r}{R}\right)\left[\int_{0}^{2\pi}\hat{\mathbf{r}}\cdot\delta\mathbf{u}|\delta\mathbf{u}|^{2}\mathrm{d}\chi\right]\mathrm{d}r.$ (6)
Here $\hat{\mathbf{r}}$ is the tangent vector at $x_{2}$ to the great circle arc passing through $x_{1},x_{2}$. A schematic of the quantities in equations (5) and (6) can be found in Fig. 1.
Figure 1: Schematic of the spherical coordinate system and elements described in equations (5) and (6).
Since the sphere is compact, it only makes sense to consider compactly-supported filter kernels. One such compactly supported filter kernel is the “standard mollifier” used by Kuzzay et al. (2015, 2017); Faranda et al. (2018), defined by:
@@
To compute the $\mathcal{D}_{L}(\mathbf{u})$ field used in this study, we first calculate velocity increments from the $(u,\upsilon,w)$ field. Then, an angular average (around a circle of great circle radius $r$) is applied at each point. The angular average is taken in the local spherical cap. The calculation of the increments is made by assuming a tangent quadratic structure. There is negligible error in the tropical regions compared to full spherical geometry, and at displacements of $\sim$40 degrees, the error in distances is $\sim$10%. This is the quantity in square brackets in (6). Our method is also less sensitive to noise because gradients are computed analytically on smooth filter kernels rather than on discrete model fields as is done in other methods (e.g. Aluie et al. (2018)). The result of the angular integral is then integrated over colorbluegreat circle separation distances $r$, weighted by the gradient of the filter kernel.
[…]
@@ 3.1 The K-Scale simulations
The CTC km-scale model domain is being run as part of the UK Met Office’s ‘K-Scale’ initiative (Jones et al. (2023), Jones et al. (In review)) which also serves as the UK contribution to the DYAMOND protocol (Stevens et al., 2019b; Takasuka et al., 2024). The DYAMOND (DYnamics of the Atmosphere Modelled On Non-hydrostatic Domains) protocol is an international collaborative project involving modelling centres across the world exploring CPMs spanning global or large domains at high resolution and their benefits. All centres modelled a common time period. We will focus on the DYAMOND summer period in this study from 2016-08-01 to 2016-09-09. The DYAMOND Summer period was chosen to match up with the NARVAL (Stevens et al., 2019a) observational campaign during these dates. 40 days was chosen to provide a month of data after any spin up and be long enough to separate questions of predictability and internal variability. The relevant tropical climate indices during DYAMOND Summer are listed in Table 1. Of particular relevance to this study is that there is no MJO during DYAMOND Summer, meaning CCEWs can be studied as their own system rather than as part of the MJO envelope (e.g. Roundy (2008); Kiladis et al. (2009)). The output data used are 3-hourly. The simulations are free-running for 40 days, i.e. with no data assimilation scheme in place like for an operational forecast. This is to allow features to grow by themselves in the model rather than be forced by the initial conditions fed in from the data assimilation scheme.
Table 1: Climate indices during the DYAMOND Summer study period. Abbreviations are the El Niño Southern Oscillation (ENSO), Indian Ocean Dipole (IOD), Madden-Julian Oscillation (MJO), and Quasi-Biennial Oscillation (QBO). The indices were calculated for ENSO using the Nino3.4 index from HadISST data (Rayner et al., 2003); the IOD using the Dipole Mode Index from HadISST data; the MJO using the Realtime Multivariate MJO index from NOAA OLR and NCEP/NCAR reanalysis winds data (Wheeler and Hendon, 2004; Kalnay et al., 1996); the QBO from NCEP/NCAR reanalysis data (Kalnay et al., 1996); and the BSISO using the all-season OLR-based MJO index (OMI, Kiladis et al. 2014).
@@
ENSO La Niña (Nino3.4r $\sim$ -1C)
IOD Negative decaying (DMI -0.4 $\to$ -0.2C)
@@
QBO Westerly ($\sim+10$ m s-1)
BSISO Phase 7 $\to$ 8 (weak) and decaying
For the analysis carried out in this study, we compare two UM (Brown et al., 2012) coupled to the Joint UK Land Environment Simulator (JULES, Best et al. 2011; Clark et al. 2011) simulations from the K-Scale hierarchy (Jones et al., In review) with differing science configurations. One simulation (referred to as GAL9 N2560 or GAL9 for short throughout) makes use of the Global Atmosphere and Land version 9 (GAL9, Walters et al. 2019) configuration. The other simulation (RAL3.2 N2560, RAL3 for short) uses the Regional Atmosphere and Land version 3.2 (RAL3.2, Bush et al. 2025). Details of the differences between the two simulations are summarised in Table 2.
Both models have identical grids and cover the CTC domain from 40S to 26N. These bounds were chosen to meet the Himalayas in the north and encompass all of southern Africa and the Indian Ocean tropical cyclone basin. The two models are both driven by a global N1280 ($\Delta x\sim 10$ km) model with GAL9 physics. The key difference between GAL9 and RAL3 being investigated in this study is the convection scheme (or lack thereof), with most other features being kept the same.
GAL9 parametrizes deep conveciton. This convection parametrization is not active in RAL3. This means that convection is permitted to explicitly evolve through the model dynamics on the grid. At this resolution there is still a lot of sub-grid convection, and this is handled by the turbulence scheme. The 3D turbulence scheme employed by RAL3 does not force artificial smoothness like the parametrization scheme in GAL, which is how explicit circulations are allowed to develop close to the grid scale. There is additional sub-grid vertical mixing in RAL3 above the boundary layer that is parametrized by a Smagorinsky-Lilly-type eddy diffusivity. Other differences between RAL3 and GAL9 that must be noted are the different cloud fraction schemes: GAL9 uses the PC2 prognostic scheme (Wilson et al., 2008), and RAL3 the bimodal diagnostic scheme (Van Weverberg et al., 2021b, a). The microphysics in GAL9 is parametrized using the Wilson and Ballard (1999) single moment scheme; and RAL3 uses the double moment Cloud AeroSol Interacting Microsphysics scheme (CASIM, Field et al. 2023). In the boundary layer, GAL9 employs the 1D turbulent mixing scheme of Lock et al. (2000), and RAL3 employs a blended 3D Smagorinsky and 1D scheme, described in Boutle et al. (2014).
@@ 3.2 Equatorial wave filtering
Equatorial wave dynamical fields are extracted from the model or analysis $(u,\upsilon,Z)$ (zonal wind, meridional wind, geopotential height) fields using the 2D-in-Space Parabolic Cylinder Function (2DS-PCF) method developed by Yang et al. (2003). The 2DS-PCF method is applied independently on individual pressure levels, and the vertical structure of the waves can then be inferred by analysis of multiple layers.
Due to the relatively short time period covered by our simulations, the “realtime” extension of the method (Yang et al., 2021) is used, whereby a long analysis period is appended to the start of the forecast to enable a time filter with a very broad window so that eastward and westward signals can be distinguished before spatial projection. There are multiple ways to identify equatorial waves from modelled and observed flows, with Knippertz et al. (2022) calling for a combination of wavenumber-frequency filtering and spatial projection methods. Since we are concerned here with kinetic energy and the 3D dynamical structure of the waves, the 2DS-PCF method is the most suitable for our purposes.
@@ 4.1 Simulated waves and precipitation
Hovmöller (longitude-time) diagrams of quantities averaged across 10S-10N are shown in Fig. 3 - the first 20 days are chosen for clarity in the figure, and to allow us to analyse the evolution of simulated features from the same initial conditions. This diagram gives us a first look at the key differences between the RAL3 and GAL9 models and the observations. In terms of precipitation: the GAL9 model has many more stationary features than the RAL3 or observations (e.g. see the features around 90W at the start of the simulation), and its eastward propagating features are quite weak. This key difference between explicit and parametrized convection in simulating propagating precipitation is also found in the study of Dias et al. (2018). The abundance of propagating precipitation in the RAL3 model is more consistent with the observations than the GAL, but appears to simulate too much rain (see for example in the Indian Ocean region around 90E). These features of the models compared to the observations are also reflected in Table 3, where GAL9 simulates a higher mean precipitation compared to GPM, albeit with a smaller variance. RAL3 simulates similar variance to the observations, but with a mean that is higher still. All differences between the datasets are significantly different at the 99% confidence level due to their large size. These differences in precipitation mean and variances between the km-scale models and observations are consistent with the findings of Judt and Rios-Berrios (2021), where they analysed the DYAMOND Summer runs from the MPAS-A model.
@@
Where the wave contours travel with the precipitation, this is a clear signal of a Convectively Coupled Equatorial Wave (CCEW). The Convectively Coupled Kelvin Wave (CCKW) initialised on 2nd August at 90E and propagating across the Maritime Continent is present in both models in Figs. 3 (b) and (c) as well as in the analysis/observations in Fig. 3 (a). In the observations, this wave propagates eastward until it encounters a convectively coupled Rossby (R1) wave and WMRG over the Maritime Continent at about 120E, resulting in a large amount of precipitation being deposited over the region, and the slowing of the CCKW. This interaction does not occur in the same way in the models in (b) and (c), with there being much weaker R1 and WMRG waves in the models; in RAL3 the KW slows much like in the observations, but the deposition of large amounts of rain is difficult to pick out amongst the very abundant rain in that model, and in the GAL9 the CCKW is terminated altogether. This interaction between the Kelvin, R1 and WMRG waves does not occur in the same way in the two models because the simulated propagation speeds of the waves in RAL3 and GAL9 are different (see Table 4). In GAL, the CCKW propagates a similar distance making it as far as the dateline in the central Pacific by Aug 11th, where it also encounters a Rossby wave, albeit in the opposite phase compared to the observations. In the GAL9 model at this point there is also not a strong WMRG like there is in the analysis. In the RAL3, the CCKW propagates farther, continuing after interaction with the easterly-phase R1 wave and terminating 180 degrees and 2 weeks later at 90W, presumably encountering the Andes. This difference may however be a factor of the free-running simulations having diverged from the analysis by 10 days out.
@@
The dominant phase speeds of the three main EW types for the UM analysis and the two models are shown in Table 4. The observed phase speeds from the analysis were cross-referenced with the extremely comprehensive analyses of Yang et al. (2007b, 2009) where they used 15 years of reanalysis data and satellite-observed convective fields, and are comparable. The models simulate a faster Kelvin wave, but the variability ranges overlap. This is the case for all the models and waves. It is also interesting to note that the models always simulate a higher standard deviation, particularly the GAL. This could be because of a reduced degree of convective coupling in GAL, allowing for a wider range of ‘dry’ (uncoupled with convection) propagating waves. Convective coupling reduces phase speed in CCEWs (Kiladis et al., 2009). For the R1 wave, the GAL9 simulates the same dominant phase speed with a higher variability, and the RAL3 a much faster phase speed with a much larger variability. The reason for this is because the distribution of phase speeds across the record calculated from the Radon transform (not shown) is bimodal. In the analysis these peaks are relatively close together at around 9 and 13 m s-1. In the RAL3, the faster peak is strongest at the 15 m s-1 value reported in the table, and the second, weaker peak, is at around 6 m s-1, hence the large variability range.
@@
The differences in the simulation of the 3 key wave types can be illustrated by a longitude-height plot of standard deviation in time of wave velocity fields as shown in Fig. 4. The six-week average structures in the UM analysis presented here are similar to the 15-year average structures presented in Yang et al. (2007a, 2009). As previously reported by these earlier studies, we expect to see the strongest wave activity in the upper-troposphere for the westward waves and around the tropopause for the Kelvin waves. This is due to the highest wave amplitude being at these levels. The peak in Kelvin wave variability around 90E at upper-tropospheric levels in the analysis is not fully simulated by either model. It is slightly weaker and further east in RAL3 (c), and not well defined in GAL9 (b). There is another peak west of 90W around 200 hPa - this is likely associated with Kelvin waves propagating across the Pacific and terminating at the Andes. This peak is stronger than analysis in GAL, and not present in RAL. The reason for this being stronger in GAL9 and not present in RAL3 could be due to the presence of the gravity wave drag (GWD) parametrization in the GAL9 simulation. This parametrization simulates the subgrid response to the flow blocking by unresolved orography. The N2560 resolution is in the ‘grey zone’ of the orographic GWD (some orography will be resolved, but only a fraction), meaning it is possible that the GWD is simulating too much drag in GAL, and without the parametrization in RAL3, not enough is simulated. Further analysis of this particular parametrization is beyond the scope of this study; see Vosper et al. (2016) for details.
@@
The R1 waves in the analysis in Fig. 4 (d) exhibit activity centred around 200 hPa in the western hemisphere (WH) - a pattern also seen in reanalysis in Yang et al. (2009). The GAL9 model simulation (e) has a less pronounced peak with wave activity spread over a wider longitudinal range, while the RAL3 (f) has an activity centre that is also weaker but more concentrated, in this case around 0. The middle and lower troposphere structure of the R1 wave in the analysis consists of a peak in activity from 90E and eastwards. For the WMRG waves in Fig. 4 (g), there are peaks in activity centred around 200 hPa mostly in the WH (also seen in the analysis product in Yang et al. (2009)). GAL9 (h) represents this quite similarly, whereas RAL3 (i) simulates a very strong peak in activity just east of 0, and the peaks in the WH are weaker than in the analysis and GAL. The differences between GAL9 and RAL3 may just be due to noise given the short record analysed, but importantly both are able to represent the general behaviour of variability in time throughout the tropical atmosphere, shown by their similarity to the UM analysis which in turn is qualitatively similar to the patterns displayed in Yang et al. (2009).